OpenStudy (anonymous):

Find the indicated limit lim->infinity (e^2x - 1)/(x)

OpenStudy (anonymous):

use L'Hopital's rule...

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

page 2, second column... http://tutorial.math.lamar.edu/pdf/Calculus_Cheat_Sheet_All.pdf

OpenStudy (sumeer):

take the derivative of (e^2x-1) and (x). it ends (2e^2x)/1

OpenStudy (anonymous):

and then what? @sumeer

OpenStudy (sumeer):

infinite i think :)

OpenStudy (anonymous):

and then what? @sumeer

OpenStudy (sumeer):

the result of question is positive infinite

OpenStudy (anonymous):

\[ L=\lim_{x\to\infty}\frac{e^{2x}-1}{x} \] and positive infinity or negative infinity?

OpenStudy (anonymous):

yes the answer for \(x\to+\infty\) is \(\infty\). he function is a continuously increasing function but, for \(x\to-\infty\), the answer would be "0"

OpenStudy (anonymous):

positive infinity

OpenStudy (anonymous):

did you study L'hopital's rules?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so, would it apply here?

OpenStudy (anonymous):

if you substitute x=infnity, you;'d get (infinity/infinity) hence, you have to use L'hopital's rule

OpenStudy (anonymous):

after applying it once, what do you get?

OpenStudy (anonymous):

umm I got lim x-> infinity e^2x - 1/x, lim x-> infinity e^ex - 1/x, lim x -> infinity d(e^2x - 1)/dx over dx/dx = e^2

OpenStudy (anonymous):

no comprehendo. use the equation editor below the reply box so the work is legible.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[ L=\lim_{x\to\infty}\frac{2e^{2x}}{1}=2\lim_{x\to\infty}e^{2x}\\ L=2\times e^{2\times\infty}\\ \boxed{L=\infty} \]

OpenStudy (anonymous):

similarly, what would happen if ... \(x\to-\infty\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok